1. Roman mixes 12 liters of 8% acid solution with a 20% acid solution, which results in a 16% acid solution. Find the number of liters of 20% acid solution in the new mixture. Construct a table to organize the information, and show all of your work to solve the problem.
Answer by stanbon(75887)

Roman mixes 12 liters of 8% acid solution with a 20% acid solution, which results in a 16% acid solution. Find the number of liters of 20% acid solution in the new mixture. Construct a table to organize the information, and show all of your work to solve the problem.
-------
Equation:
acid + acid = acid
0.08*12 + 0.20x = 0.16(12+x)
-------
8*12 + 20x = 16*12 + 16x
4x = 8*12
x = 24 Liters (amt. of 20% solution needed)
--------------------------------------------
Cheers
Stan H.
--------------------------------